# Feng-Yu Wanga), Chenggui Yuanb), Xiao-Yu Zhaoa)

### v1

@@ 1 Introduction
The dimension-free Harnack inequality with power was first introduced in [13] to study the log-Sobolev inequality on Riemanian manifolds, and has been intensively extended and applied to derive regularity estimates many for SDEs (stochastic differential equations), SPDEs (stochastic partial differential equations), path dependent SDEs, and distribution-dependent SDEs, see [16, 12] and references therein. As a limit version when the power goes to infinite, the log-Harnack inequality has been introduced in [14] to characterize the curvature lower bounded and entropy-cost estimates, and has been extended to metric measure spaces [1, 10]. When the noise of a stochastic system is weak enough such that the log-Harnack inequality is not available, the asymptotic log-Harnack inequality has been studied in [5, 20] for path-dependent SDEs and SPDEs, which in particular implies an asymptotic gradient estimate.
@@
This is a Polish space under the norm $\|\cdot\|_{\tau}$. Let $\mathscr{P}$ and $\mathscr{P}(\mathscr{C}_{\tau})$ be the set of all probability measures on $(\mathbb{R}^{d},\mathscr{B}(\mathbb{R}^{d}))$ and $(\mathscr{C}_{\tau},\mathscr{B}(\mathscr{C}_{\tau}))$, respectively, equipped with the weak topology. Let $\mathscr{B}_{b}(\mathscr{C}_{\tau})$ be the class of bounded measurable functions on $\mathscr{C}_{\tau}$, and $\mathscr{B}_{b}^{+}(\mathscr{C}_{\tau})$ the set of strictly positive functions in $\mathscr{B}_{b}(\mathscr{C}_{\tau})$.
@@
$\text{\rm{d}}X(t)=b(t,X_{t},\mathscr{L}_{X_{t}})\text{\rm{d}}t+\sigma(t,X(t))\mathrm{d}W(t),\quad t\geq 0,$ (1.2)
@@
$b:\mathbb{R}^{+}\times\mathscr{C}_{\tau}\times\mathscr{P}(\mathscr{C}_{\tau})\rightarrow\mathbb{R}^{d},\ \ \ \sigma:\mathbb{R}^{+}\times\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}\otimes\mathbb{R}^{d}$
@@
$\mathscr{P}_{k}(\mathscr{C}_{\tau}):=\big{\{}\mu\in\mathscr{P}(\mathscr{C}_{\tau}):\|\mu\|_{k}:=\mu(|\cdot|^{k})^{\frac{1}{k}}<\infty\big{\}},$
where for $k=0$ we set $\mu(|\cdot|^{k})^{\frac{1}{k}}=1$ such that $\mathscr{P}_{0}(\mathscr{C}_{\tau})=\mathscr{P}(\mathscr{C}_{\tau}).$ When $k>0$, $\mathscr{P}_{k}(\mathscr{C}_{\tau})$ is a Polish space under the $L^{k}$-Wasserstein distance,
@@
where $|\mu-\nu|$ is the total variation of $\mu-\nu$. According to [12, Remark 3.2.1], for any $k>0$, there exists a constant $c>0$ such that
@@
    $\int_{0}^{t}\mathbb{E}\big{[}|b(r,X_{r},\mathscr{L}_{X_{r}})|+\|\sigma(r,X(r))\|^{2}\big{|}\mathscr{F}_{0}\big{]}\mathrm{d}r<\infty,\ \ t\geq 0,$
@@
    $X(t)=X(0)+\int_{0}^{t}b(r,X_{r},\mathscr{L}_{X_{r}})\mathrm{d}r+\int_{0}^{t}\sigma(r,X(r))\mathrm{d}W(r),\ \ t\geq 0.$
@@
    The SDE (1.2) is called well-posed for distributions in $\tilde{\mathscr{P}}_{k}(\mathscr{C}_{\tau})$, if it is both strongly and weakly well-posed for distributions in $\tilde{\mathscr{P}}_{k}(\mathscr{C}_{\tau})$. In this case, for any $\gamma\in\tilde{\mathscr{P}}_{k}(\mathscr{C}_{\tau}),$ let $P_{t}^{*}\gamma=\mathscr{L}_{X_{t}^{\gamma}}$ with $L_{X_{0}^{\gamma}}=\gamma$.
[…]
$P_{t}f(\xi):=\mathbb{E}[f(X_{t}^{\xi})],\ \ t\geq 0,\ f\in\mathscr{B}_{b}(\mathscr{C}_{\tau}).$
@@
2 Main Results
Let $k\geq 0$. To ensure the well-posedness of solutions to (1.2) with singular drift for distributions in $\mathscr{P}_{k}(\mathscr{C}_{\tau})$, we decompose $b$ as
and impose the following assumptions on $b^{(0)},b^{(1)}$ and $\sigma$. For any $\xi\in\mathscr{C}_{\tau}$, let $\xi^{0}\in\mathscr{C}_{\tau}$ be defined as
@@ 2 Main Results
  • ($H_{2}$)
    There exists a function $H\in L^{1}_{loc}([0,\infty);(0,\infty))$ and constants $K>0$, $\alpha\in[0,1]$ such that
    $|b^{(1)}(t,\xi,\mu)-b^{(1)}(t,\eta,\nu)|^{2}\leq{K}\|\xi-\eta\|_{\tau}^{2}+H(t)\mathbb{W}_{k,var}(\mu,\nu)^{2},$
    $|b^{(1)}(t,\xi,\mu){-b^{(1)}(t,\xi^{0},\mu)}|\leq{K}\left\{1+\|\xi\|_{\tau}^{\alpha}+\|\mu\|_{k}\right\},$ (2.1)
@@
  • ($H_{3}$)
    There exist constants $\{(p_{i},q_{i})\}_{0\leq i\leq l}\in\mathcal{K}$ with $l\geq 1$, $p_{i}>2$, and functions $0\leq f_{i}\in\cap_{n\in\mathbb{N}}\tilde{L}^{p_{i}}_{q_{i}}(n)$, $0\leq i\leq l$, such that
    $|b^{(0)}|\leq f_{0},\ \ \|\nabla\sigma\|\leq\sum_{i=0}^{l}{f_{i}}.$
@@
To establish the asymptotic log-Harnack inequality, we assume that the following monotone assumption holds, as in [5] in the distribution-free case. An important application of this inequality is the gradient estimate. For any function $f$ on $\mathscr{C}_{\tau}$, let
be the Lipschitz constant of a function $f$ at point $\xi\in\mathscr{C}_{\tau}$, which coincides with the norm of the gradient $\nabla f(\xi)$ if $f$ is Gâdeux differentiable at $\xi$. We denote $f\in C_{b,L}(\mathscr{C}_{\tau})$ if
$\|f\|_{\infty}+\|\nabla f\|_{\infty}<\infty.$
@@ 1. $(H^{\prime})$
    For any $t\geq 0$ and $\mu\in\mathscr{P}(\mathscr{C}_{\tau})$, $b(t,\cdot,\mu)\in C(\mathscr{C}_{\tau};\mathbb{R}^{d})$, and there exists a constant $K>0$ such that $\sup_{t\geq 0}|b(t,{\bf 0},\delta_{\bf 0})|\leq{K}$ and
    $\langle b(t,\xi,\mu)-b(t,\eta,\nu),\xi(0)-\eta(0)\rangle^{+}+\|\sigma(t,\xi(0))-\sigma(t,\eta(0))\|^{2}$
    $\leq K\big{(}\|\xi-\eta\|_{\tau}^{2}+\mathbb{W}_{2}(\mu,\nu)^{2}\big{)},\ \ t\geq 0,\ \xi,\eta\in\mathscr{C}_{\tau},\ \mu,\nu\in\mathscr{P}_{2}(\mathscr{C}_{\tau}).$
@@ 2. $(2)$
    $P_{t}\log f(\nu)\leq$ $\log P_{t}f(\mu)+{c\text{\rm{e}}^{ct}}\mathbb{W}_{2}(\mu,\nu)^{2}+c\text{\rm{e}}^{-\tau_{0}t}\|\nabla\log f\|_{\infty}\mathbb{W}_{2}(\mu,\nu)$ (2.6)
@@
    $|P_{t}f(\mu)-P_{t}f(\nu)|\leq\mathbb{W}_{2}(\mu,\nu)\big{[}c\text{\rm{e}}^{-\tau_{0}t}\|\nabla f\|_{\infty}+{2\sqrt{c\text{\rm{e}}^{ct}}}\|f\|_{\infty}\big{]}.$ (2.7)
3 Path Dependent SDEs with Infinite Memory
In this section, we consider the following path-distribution independent SDE with infinite memory on $\mathbb{R}^{d}$:
3.1 Well-posedness and Lipschitz continuity in initial value
and make the following assumptions on $b^{(0)},b^{(1)}$ and $\sigma$.
  • $(A_{1})$
    $b^{(0)}$ and $a:=\sigma\sigma^{*}$ satisfy $(H_{1})$ and $(H_{3})$.
  • $(A_{2})$
    For every $n>0$, there exists a constant $K_{n}>0$ such that
    Moreover, there exists a constant $K>0$ such that
    $|b^{(1)}(t,\xi)|\leq K(1+\|\xi\|_{\tau}),\ \ {t\geq 0},\xi\in\mathscr{C}_{\tau}.$ (3.3)
Theorem 3.1.
Assume $(A_{1})$ and $(A_{2})$. Then (3.1) is well-posed for any initial value in $\mathscr{C}_{\tau}$, and for any constants $k,T>0$, there exists a constant $c>0$, such that
Let $X_{t}^{\xi}$ be the segment solution with $X_{0}^{\xi}=\xi$. To ensure the Lipschiz continuity of $X_{t}^{\xi}$ in $\xi$, we strengthen $(A_{2})$ to
  • $(A_{2}^{\prime})$
    There exist constants $K>0$ and $\alpha\in[0,1]$ such that
    $|b^{(1)}(t,\xi)-b^{(1)}(t,\eta)|\leq K\|\xi-\eta\|_{\tau},$
    $|b^{(1)}(t,\xi)-b^{(1)}(t,\xi^{0})|\leq K\left(1+\|\xi\|_{\tau}^{\alpha}\right),\ \ t\in[0,T],\ \xi,\eta\in\mathscr{C}_{\tau}.$
Theorem 3.2.
Assume $(A_{1})$ and $(A_{2}^{\prime})$. Then for any constants $\varepsilon\in(0,1)$ and $k,T\geq 1$, there exists a constant $c>0$ such that
To prove Theorem 3.1 and Theorem 3.2, we present some lemmas below.
@@ 3.1 Well-posedness and Lipschitz continuity in initial value
Assume $(A_{1})$, $(A_{2}^{\prime})$ and $\|b^{(1)}\|_{\infty}<\infty$. Then for any initial value $\xi\in\mathscr{C}_{\tau}$, (3.1) has a unique non-explosive strong solution satisfying (3.4), and for any $T\in(0,\infty)$ there exists a constant $c>0$ such that
@@
The desired estimate follows from the proof of [23, Lemma 4.1] with $\|\cdot\|_{\tau}$ in place of uniform norm $\|\cdot\|_{\mathscr{C}}$ on the path space with finite time interval. ∎
@@
where $|B(0,r)|$ is the volume of $B(0,r):=\{y:|y|<r\}$. The following result is taken from [19, Lemma 2.1].
@@ 2. $(2)$
Assume $(A_{1})$ and $(A_{2}^{\prime})$. For any $T\in(1,\infty)$, there exist constants $\beta_{T},k_{T}>0$ such that for any solution $(X_{t})_{t\in[0,T]}$ to (3.1),
@@
By [18, Theorem 3.2] and [22, Theorem 2.1], there exists a constant $\lambda_{0}>0$ such that for any $\lambda\geq\lambda_{0}$, the PDE for $u(t,\cdot):\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}$
@@
$\|u\|_{\infty}+\|\nabla u\|_{\infty}+\|\nabla^{2}u\|_{\tilde{L}_{q_{0}}^{p_{0}}(T)}\leq\frac{1}{2},$ (3.14)
@@
For any $t\in[0,T]$, let $Y(t)=\Theta(t,X(t))$. By Itô’s formula and (3.12), we obtain
@@
Combining this with $(A_{1})$, the second inequality in $(A_{2}^{\prime})$, (3.14) and Itô’s formula, we find a constant $c_{0}>0$ such that
@@
$l(t):=\text{\rm{e}}^{2\alpha\tau t}\{|Y(t)|^{2}+1\}^{\alpha}+c_{1}\left(1+\|Y_{0}\|_{\tau}^{2\alpha}\right),$
$\bar{l}_{t}:=\sup_{s\in[0,t]}l(s)\ \ t\in[0,T],$
@@
Assume $(A_{1})$ and $(A_{2}^{\prime})$. For any constants $T,p,q>1$ such that $(2p,2q)\in\mathscr{K}$ and $\varepsilon>0$, there exist constants $c=c(T,p,q,\varepsilon)>0$ and $\theta=\theta(T,p,q)>0$ such that the solution to (3.1) satisfies
@@
By (3.22), $(A_{1})$, the second inequality in $(A_{2}^{\prime})$, and applying Krylov’s estimate in [22, Theorem 3.1] for $\bar{b}(t,x)$ in place of $b(t,x)$, see also [12, Theorem 1.2.3(2)], we find constants $k_{2}$ depending on $\varepsilon$ and $k_{3}>0$ independent of $\varepsilon$ such that
$\mathbb{E}\left[\int_{0}^{t\wedge\tau_{n}}|f_{s}|\left((X(s)\right)\text{\rm{d}}s\bigg{|}\mathscr{F}_{0}\right]$ (3.24)
$\leq$ $\left\{k_{2}+\varepsilon\left[\mathbb{E}\left(\int_{0}^{t\wedge\tau_{n}}|\hat{b}(s,X_{s}))|^{2}\text{\rm{d}}s\bigg{|}\mathscr{F}_{0}\right)\right]^{\frac{1}{2}}\right\}\|f\|_{\tilde{L}_{q}^{p}(0,t)}$
@@
By applying the Krylov estimate in [12, Theorem 1.2.4], we find constant $k>0$ independent of $s$ such that
$\mathbb{E}_{\mathbb{Q}_{n}}\left[\exp\left(\lambda\int_{0}^{t\land\tau_{n}}|f_{r}|\left(X(r)\right)\text{\rm{d}}r\right)\bigg{|}\mathscr{F}_{0}\right]\leq\exp\Big{[}k+k\|f\|_{\tilde{L}_{q}^{p}(0,t)}^{k}\Big{]},\ \ t\in[0,T].$ (3.27)
On the other hand, for constants $K$ in $(A_{1})$ and $(A_{2}^{\prime})$, and $\beta_{T}$ in Lemma 3.5, let
@@
Thus, taking $\tilde{p}\in[1,p_{0}]$ with $\tilde{p}-1$ small enough such that ${k_{T}}k(\tilde{p})\leq 2\varepsilon$, by $(A_{1})$, $(A_{2}^{\prime})$ and (3.10), we find a constant $\beta=\beta(p,q,\varepsilon)>0$ such that

### v2

@@ 1 Introduction
The dimension-free Harnack inequality with power was first introduced in [11] to study the log-Sobolev inequality on Riemanian manifolds, and has been intensively extended and applied to derive regularity estimates many for SDEs (stochastic differential equations), SPDEs (stochastic partial differential equations), path dependent SDEs, and distribution-dependent SDEs, see [13, 10, 9] and references therein. As a limit version when the power goes to infinite, the log-Harnack inequality has been introduced in [12] to characterize the curvature lower bounded and entropy-cost estimates, and has been extended to metric measure spaces [1, 8]. When the noise of a stochastic system is weak such that the log-Harnack inequality is not available, the asymptotic log-Harnack inequality has been studied in [5, 16] for path-dependent SDEs and SPDEs, which in particular implies an asymptotic gradient estimate.
@@
It is well known that $(C((-\infty,0];\mathbb{R}^{d}),\|\cdot\|_{\infty})$ is complete but not separable, so is $(\mathscr{C}_{\tau},\|\cdot\|_{\tau})$ due to the isometric
$\mathscr{C}_{\tau}\ni\xi:=(\xi_{s})_{s\in(-\infty,0]}\mapsto\text{\rm{e}}^{\tau\cdot}\xi:=(\text{\rm{e}}^{\tau s}\xi_{s})_{s\in(-\infty,0]}\in C((-\infty,0];\mathbb{R}^{d}).$
Let $\mathscr{P}$ and $\mathscr{P}(\mathscr{C}_{\tau})$ be the set of all probability measures on $(\mathbb{R}^{d},\mathscr{B}(\mathbb{R}^{d}))$ and $(\mathscr{C}_{\tau},\mathscr{B}(\mathscr{C}_{\tau}))$, respectively, equipped with the weak topology. Let $\mathscr{B}_{b}(\mathscr{C}_{\tau})$ be the class of bounded measurable functions on $\mathscr{C}_{\tau}$, and $\mathscr{B}_{b}^{+}(\mathscr{C}_{\tau})$ the set of strictly positive functions in $\mathscr{B}_{b}(\mathscr{C}_{\tau})$.
@@
$\text{\rm{d}}X(t)=b(t,X_{t},\mathscr{L}_{X_{t}})\text{\rm{d}}t+\sigma(t,X_{t})\mathrm{d}W(t),\quad t\geq 0,$ (1.2)
@@
$b:\mathbb{R}^{+}\times\mathscr{C}_{\tau}\times\mathscr{P}(\mathscr{C}_{\tau})\rightarrow\mathbb{R}^{d},\ \ \ \sigma:\mathbb{R}^{+}\times\mathscr{C}_{\tau}\rightarrow\mathbb{R}^{d}\otimes\mathbb{R}^{d}$
@@
$\mathscr{P}_{k}(\mathscr{C}_{\tau}):=\big{\{}\mu\in\mathscr{P}(\mathscr{C}_{\tau}):\|\mu\|_{k}:=\mu(\|\cdot\|_{\tau}^{k})^{\frac{1}{k}}<\infty\big{\}},$
where for $k=0$ we set $\mu(\|\cdot\|_{\tau}^{k})^{\frac{1}{k}}=1$ such that $\mathscr{P}_{0}(\mathscr{C}_{\tau})=\mathscr{P}(\mathscr{C}_{\tau}).$ When $k>0$, $\mathscr{P}_{k}(\mathscr{C}_{\tau})$ is a Polish space under the $L^{k}$-Wasserstein distance,
@@
where $|\mu-\nu|$ is the total variation of $\mu-\nu$. According to [10, Remark 3.2.1], for any $k>0$, there exists a constant $c>0$ such that
@@
    $\int_{0}^{t}\mathbb{E}\big{[}|b(r,X_{r},\mathscr{L}_{X_{r}})|+\|\sigma(r,X_{r})\|^{2}\big{|}\mathscr{F}_{0}\big{]}\,\mathrm{d}r<\infty,\ \ t\geq 0,$
@@
    $X(t)=X(0)+\int_{0}^{t}b(r,X_{r},\mathscr{L}_{X_{r}})\,\mathrm{d}r+\int_{0}^{t}\sigma(r,X_{r})\,\mathrm{d}W(r),\ \ t\geq 0.$
@@
    The SDE (1.2) is called well-posed for distributions in $\tilde{\mathscr{P}}_{k}(\mathscr{C}_{\tau})$, if it is both strongly and weakly well-posed for distributions in $\tilde{\mathscr{P}}_{k}(\mathscr{C}_{\tau})$. In this case, for any $\xi\in\mathscr{C}_{\tau}$ such that $\mathscr{L}_{\xi}\in\tilde{\mathscr{P}}_{k}(\mathscr{C}_{\tau})$, let
    $P_{t}^{*}\gamma=\mathscr{L}_{X_{t}^{\xi}},\ \ \gamma=\mathscr{L}_{\xi}$
    and denote
    $P_{t}f(\gamma):=\mathbb{E}[f(X_{t}^{\xi})]=\int_{\mathscr{C}_{\tau}}f\,\text{\rm{d}}\left\{P_{t}^{*}\gamma\right\},\ \ \gamma=\mathscr{L}_{\xi},\,t\geq 0,\,f\in\mathscr{B}_{b}(\mathscr{C}_{\tau}),$
    where $(X_{t}^{\xi})_{t\geq 0}$ is the unique solution to (1.2) with $X_{0}^{\xi}=\xi$.
@@
2 The Singular Case: Well-Posedness and Lipschitz Continuity in Initial Value
In this section, we let $k\geq 0$ and consider (1.2) with singular drifts and $\sigma(t,\xi)=\sigma(t,\xi(0))$.
2.1 Path Dependent SDEs with Infinite Memory
In this part, we consider the following path dependent SDE with infinite memory on $\mathbb{R}^{d}$:
and make the following assumptions on $b^{(0)},b^{(1)}$ and $\sigma$. For any $\xi\in\mathscr{C}_{\tau}$, let $\xi^{0}\in\mathscr{C}_{\tau}$ be defined as
@@ 2.1 Path Dependent SDEs with Infinite Memory
  2. ($A_{2}$)
    There exist constants $\{(p_{i},q_{i})\}_{0\leq i\leq l}\in\mathscr{K}$ with $l\geq 1$, $p_{i}>2$, and functions $0\leq f_{i}\in\cap_{n\in\mathbb{N}}\tilde{L}^{p_{i}}_{q_{i}}(n)$, $0\leq i\leq l$, such that
    $|b^{(0)}|\leq f_{0},\ \ \|\nabla\sigma\|\leq\sum_{i=0}^{l}{f_{i}}.$
  3. $(A_{3})$
    For every $n>0$, there exists a constant $K_{n}>0$ such that
    Moreover, there exists a constant $K>0$ such that
    $|b^{(1)}(t,\xi)-b^{(1)}(t,\xi^{0})|\leq K(1+\|\xi\|_{\tau}),\ \ {t\geq 0},\xi\in\mathscr{C}_{\tau}.$ (2.3)
Theorem 2.1.
Assume $(A_{1})$$(A_{3})$. Then (2.1) is well-posed for any initial value in $\mathscr{C}_{\tau}$, and for any constants $k,T>0$, there exists a constant $c>0$, such that
Let $X_{t}^{\xi}$ be the segment solution with $X_{0}^{\xi}=\xi$. To ensure the Lipschitz continuity of $X_{t}^{\xi}$ in $\xi$, we strengthen $(A_{3})$ to
  • $(A_{3}^{\prime})$
    $\sup_{t\geq 0}|b^{(1)}(t,{\bf 0})|<\infty$, and there exist constants $K>0$ and $\alpha\in[0,1]$ such that
    $|b^{(1)}(t,\xi)-b^{(1)}(t,\eta)|\leq K\|\xi-\eta\|_{\tau},$
    $|b^{(1)}(t,\xi)-b^{(1)}(t,\xi^{0})|\leq K\left(1+\|\xi\|_{\tau}^{\alpha}\right),\ \ t\geq 0,\ \xi,\eta\in\mathscr{C}_{\tau}.$
Theorem 2.2.
Assume $(A_{1})$, $(A_{2})$ and $(A_{3}^{\prime})$. Then for any constants $\varepsilon\in(0,1)$ and $k,T\geq 1$, there exists a constant $c>0$ such that
2.2 Path-Distribution Dependent SDEs with Infinite Memory
To ensure the well-posedness of solutions to (1.2) with singular drift for distributions in $\mathscr{P}_{k}(\mathscr{C}_{\tau})$, we decompose $b$ as
and impose the following assumptions on $b^{(0)},b^{(1)}$ and $\sigma$.
  1. $(H_{1})$
    $b^{(0)}$ and $a:=\sigma\sigma^{*}$ satisfy $(A_{1})$ and $(A_{2})$.
  2. ($H_{2}$)
    $\sup_{t\geq 0}|b^{(1)}(t,{\bf 0},{\delta}_{\bf 0})|<\infty$, and there exists a function $H\in L^{1}_{loc}([0,\infty);(0,\infty))$ and constants $K>0$, $\alpha\in[0,1]$ such that
    $|b^{(1)}(t,\xi,\mu)-b^{(1)}(t,\eta,\nu)|^{2}\leq{K}\|\xi-\eta\|_{\tau}^{2}+H(t)\mathbb{W}_{k,var}(\mu,\nu)^{2},$ (2.6)
    $|b^{(1)}(t,\xi,\mu){-b^{(1)}(t,\xi^{0},\mu)}|\leq{K}\left\{1+\|\xi\|_{\tau}^{\alpha}+\|\mu\|_{k}\right\},$
@@ 1. $(H_{1})$
3 The Monotone Case: Well-Posedness and Asymptotic log-Harnack inequality
In this section, we consider (1.2) with monotone coefficients and establish the well-posedness and asymptotic log-Harnack inequality.
3.1 Path Dependent SDEs with Infinite Memory
Note that when the SDE
$\text{\rm{d}}X(t)=b(t,X_{t})\mathrm{d}t+\sigma(t,X_{t})\mathrm{d}W(t),\ \ t>0,\ \ X_{0}=\xi\in\mathscr{C}_{\tau}.$ (3.1)
is time-homogenous and monotone, the asymptotic log-Harnack inequality has been derived in [5, Theorem 3.1] for the Markov semigroup
@@ 3.1 Path Dependent SDEs with Infinite Memory
An important application of this inequality is the gradient estimate. For any function $f$ on $\mathscr{C}_{\tau}$, let
be the Lipschitz constant of a function $f$ at point $\xi\in\mathscr{C}_{\tau}$, which coincides with the norm of the gradient $\nabla f(\xi)$ if $f$ is Gâteaux differentiable at $\xi$. We denote $f\in C_{b,L}(\mathscr{C}_{\tau})$ if
$\|f\|_{\infty}+\|\nabla f\|_{\infty}<\infty.$
  1. $(B)$
    where ${\bf 0}\in\mathscr{C}_{\tau}$ with ${\bf 0}(r)=0,r\leq 0.$
Theorem 3.1.
[…]
$P_{t}\log f(\eta)\leq\log P_{t}f(\xi)+c\|\xi-\eta\|_{\tau}^{2}+c\text{\rm{e}}^{-\tau_{0}t}\|\nabla\log f\|_{\infty}\|\xi-\eta\|_{\tau}.$ (3.2)
Corollary 3.2.
In the situation of Theorem 3.1, the following assertions hold.
  1. $(1)$
  2. $(2)$
  3. $(3)$
    $\delta(\xi,A):=\liminf_{t\rightarrow\infty}P_{t}(\xi,A)>0.$
3.2 Path-Distribution Dependent SDEs with Infinite Memory
we assume that the following monotone assumption holds, as in [5] in the distribution-free case.
@@ 1. $(H^{\prime})$
    $\sup_{t\geq 0}|b(t,{\bf 0},\delta_{\bf 0})|<\infty$, and $b(t,\cdot,\mu)\in C(\mathscr{C}_{\tau};\mathbb{R}^{d})$ holds for any $t\geq 0$, $\mu\in\mathscr{P}(\mathscr{C}_{\tau})$. Moreover, there exist constants $K_{1},K_{2}>0$ such that and
    $\langle b(t,\xi,\mu)-b(t,\eta,\nu),\xi(0)-\eta(0)\rangle^{+}+\|\sigma(t,\xi)-\sigma(t,\eta)\|^{2}$
    $\leq K_{1}\|\xi-\eta\|_{\tau}^{2}+K_{2}\mathbb{W}_{2}(\mu,\nu)^{2},\ \ t\geq 0,\ \xi,\eta\in\mathscr{C}_{\tau},\ \mu,\nu\in\mathscr{P}_{2}(\mathscr{C}_{\tau}).$
@@ 2. $(2)$
    $P_{t}\log f(\nu)\leq$ $\log P_{t}f(\mu)+{cK_{2}\text{\rm{e}}^{ct}}\mathbb{W}_{2}(\mu,\nu)^{2}+c\text{\rm{e}}^{-\tau_{0}t}\|\nabla\log f\|_{\infty}\mathbb{W}_{2}(\mu,\nu)$ (3.5)
@@
    $|P_{t}f(\mu)-P_{t}f(\nu)|\leq\mathbb{W}_{2}(\mu,\nu)\big{[}c\text{\rm{e}}^{-\tau_{0}t}\|\nabla f\|_{\infty}+{2\sqrt{cK_{2}\text{\rm{e}}^{ct}}}\|f\|_{\infty}\big{]}.$ (3.6)
Remark 3.4.
Note that (3.5) implies (3.1) for $\mu=\delta_{\xi}$, $\nu=\delta_{\eta}$ and $K_{2}=0$.
4 Proof of Theorem 2.1 Theorem 2.2
To prove these two theorems , we present some lemmas below.
@@
Assume $(A_{1})$, $(A_{2})$, $(A_{3}^{\prime})$ and $\|b^{(1)}\|_{\infty}<\infty$. Then for any initial value $\xi\in\mathscr{C}_{\tau}$, (2.1) has a unique non-explosive strong solution satisfying (2.4), and for any $T\in(0,\infty)$ there exists a constant $c>0$ such that
@@
The desired estimate follows from the proof of [18, Lemma 4.1] with $\|\cdot\|_{\tau}$ in place of uniform norm $\|\cdot\|_{\mathscr{C}}$ on the path space with finite time interval. ∎
@@
where $|B(0,r)|$ is the volume of $B(0,r):=\{y:|y|<r\}$. The following result is taken from [15, Lemma 2.1].
@@
Assume $(A_{1})$, $(A_{2})$ and $(A_{3}^{\prime})$. For any $T\in(1,\infty)$, there exist constants $\beta_{T},k_{T}>0$ such that for any solution $(X_{t})_{t\in[0,T]}$ to (2.1),
@@
By [14, Theorem 3.2] and [17, Theorem 2.1], there exists a constant $\lambda_{0}>0$ such that for any $\lambda\geq\lambda_{0}$, the PDE for $u(t,\cdot):\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}$
@@
$\|u\|_{\infty}+\|\nabla u\|_{\infty}\leq\frac{1}{2},$ (4.9)
@@
For any $t\in[0,T]$, let $Y(t)=\Theta(t,X(t))$. By Itô’s formula in [10, Theorem 1.2.3(3)] and (4.7), we obtain
@@
Combining this with $(A_{1})$, $(A_{2})$, the second inequality in $(A_{3}^{\prime})$, (4.9) and Itô’s formula, we find a constant $c_{0}>0$ such that
@@
$l(t)=\text{\rm{e}}^{2\alpha\tau t}\{|Y(t)|^{2}+1\}^{\alpha}+c_{1}\left(1+\|Y_{0}\|_{\tau}^{2\alpha}\right),$
$\bar{l}_{t}=\sup_{s\in[0,t]}l(s)\ \ t\in[0,T],$
@@
Assume $(A_{1})$, $(A_{2})$ and $(A_{3}^{\prime})$. For any constants $T,p,q>1$ such that $(2p,2q)\in\mathscr{K}$ and $\varepsilon>0$, there exist constants $c=c(T,p,q,\varepsilon)>0$ and $\theta=\theta(T,p,q)>0$ such that the solution to (2.1) satisfies
@@
By (4.17), $(A_{1})$, $(A_{2})$, the second inequality in $(A_{3}^{\prime})$, and applying Krylov’s estimate in [17, Theorem 3.1] for $\bar{b}(t,x)$ in place of $b(t,x)$, see also [10, Theorem 1.2.3(2)], we find constants $k_{2}$ depending on $\varepsilon$ and $k_{3}>0$ independent of $\varepsilon$ such that
$\mathbb{E}\left[\int_{0}^{t\wedge\tau_{n}}|f_{s}|\left(X(s)\right)\,\text{\rm{d}}s\bigg{|}\mathscr{F}_{0}\right]$ (4.19)
$\leq$ $\left\{k_{2}+\varepsilon\left[\mathbb{E}\left(\int_{0}^{t\wedge\tau_{n}}|\hat{b}(s,X_{s})|^{2}\,\text{\rm{d}}s\bigg{|}\mathscr{F}_{0}\right)\right]^{\frac{1}{2}}\right\}\|f\|_{\tilde{L}_{q}^{p}(0,t)}$
@@
Since the Krylov estimate in [10, Theorem 1.2.4] holds for (4.21), letting $f^{n}_{r}(x)=f_{r}(x){\bf 1}_{\{|x|<n\}}$, by the argument in the proof of [18, Lemma 4.4], we find a constant $k>0$ independent of $s$ such that for any $t\in[0,T]$,
$\mathbb{E}_{\mathbb{Q}_{n}}\left[\exp\left(\lambda\int_{0}^{t}\left|f^{n}_{r}\left(X(r\wedge\tau_{n})\right)\right|\,\text{\rm{d}}r\right)\bigg{|}\mathscr{F}_{0}\right]\leq\exp\Big{[}k+k\|f\|_{\tilde{L}_{q}^{p}(0,t)}^{k}\Big{]}.$ (4.22)
On the other hand, for constants $K$ in $(A_{1})$ $(A_{2})$, $(A_{3}^{\prime})$, and $\beta_{T}$ in Lemma 4.3, let
@@
Thus, taking $\tilde{p}\in[1,p_{0}]$ with $\tilde{p}-1$ small enough such that ${k_{T}}k(\tilde{p})\leq 2\varepsilon$, by $(A_{1})$ $(A_{2})$, $(A_{3}^{\prime})$ and (4.5), we find a constant $\beta=\beta(p,q,\varepsilon)>0$ such that